solve 2 questions sow work
1. (4x)^2 + 8x^2
2. 6^2 * 6^5

Answer by lwsshak3(11628)

solve 2 questions show work
1. (4x)^2 + 8x^2
=16x^2+8x^2
=24x^2
..
2. 6^2 * 6^5=6^7=279936